OpenStudy (gabylovesyou):

Solve V = 1/3BH for B B = 3V/H B = 3H/V B = V/3H B = H/3V

OpenStudy (jhannybean):

\[V=\frac{ 1 }{ 3 }BH\] \[3V=BH\]\[\frac{ 3V }{ H }=B\]

OpenStudy (anonymous):

B = 3V/H

OpenStudy (gabylovesyou):

but how did u get it ?

OpenStudy (anonymous):

multiply by 3 on both sides and divide by h on both sides

OpenStudy (jhannybean):

first we multiply 3 to the other side,so we get \[3*V=(\frac{ 1 }{ 3 }BH)*3\] we get \[3V=BH\] then we divide by H to eliminate the H from the B\[\frac{ 3V }{ H }=BH*\frac{ 1 }{ H }\]\[\frac{ 3V }{ H }=B\]

OpenStudy (gabylovesyou):

Thank you :)

OpenStudy (jhannybean):

No Problem :)